If the area of the rectangle is 105 cm2. Its length is (4x – 5) cm and breadth is (2x – 5) cm, find the perimeter?

Answer by TimothyLamb(4379)

---
L = 4x – 5
W = 2x – 5
---
(4x – 5)(2x – 5) = 105
8xx – 20x - 10x + 25 = 105
8xx – 30x - 80 = 0
---
the above quadratic equation is in standard form, with a=8, b=-30 and c=-80
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
8 -30 -80
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 5.55136029
x = -1.80136029
---
x = 13.6509717
L = 4x - 5
L = 4*13.6509717 - 5
L = 49.6038868
W = 2x - 5
W = 2*13.6509717 - 5
W = 22.3019434
---
x = -1.80136029
L = 4x - 5
L = 4*-1.80136029 - 5
L = -12.20544116
W = 2x - 5
W = 2*-1.80136029 - 5
W = -8.60272058
---
the negative root doesn't fit the problem statement, so use the positive root:
---
answer:
p = 2L + 2W
p = 2*49.6038868 + 2*22.3019434
p = 143.8116604 cm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---